OpenStudy (anonymous):

Solve the equation: 1.4x - 2.8 = 4.2 5 10 0.5 7

OpenStudy (anonymous):

1.4x -2.8 = 4.2 1.4x = 4.2 + 2.8 x = (4.2 + 2.8)/1.4